OpenStudy (flygirlmandi):

86 is 45% of what number? A-86=0.45n B-n=0.45 x 86 C-0.45=n x 86 D-86=45n thanks :)

OpenStudy (agl202):

What do u think?

OpenStudy (safa102):

its D for sure

OpenStudy (dayakar):

let the number is n

OpenStudy (agl202):

Plz don't give direct answers out. Support the user with explanation.

OpenStudy (dayakar):

what is the 45% of n?

OpenStudy (safa102):

because u always have to follow the question and make that into your answer, so its D because 86 does = 45 percent than n the number we are trying to find out!

OpenStudy (safa102):

so its exactly like how the question say's. And D is your answer :)

OpenStudy (safa102):

can i have a medal plz!

OpenStudy (safa102):

thx! :)

OpenStudy (dayakar):

@Safa102 , incorrect answer

OpenStudy (flygirlmandi):

is the answer d or no

OpenStudy (dayakar):

it is wrong

OpenStudy (flygirlmandi):

so im not good at this stuff

OpenStudy (dayakar):

try to do it follow the instructions

OpenStudy (flygirlmandi):

wait so if 45 is considerably close to 50 which is 1/2 does that mean i have to multiply by 2 to get around the right answer ?

OpenStudy (dayakar):

% means per cent or per 100 % = 1/100 15% = 15/100 = 0.15 is it clear

OpenStudy (dayakar):

45% = ?

OpenStudy (flygirlmandi):

.45

OpenStudy (dayakar):

compare with 15% which i did above

OpenStudy (dayakar):

sorry , ur right

OpenStudy (flygirlmandi):

so 45 is 0.45 but i dont see why that helps (sorry if i seem rude)

OpenStudy (dayakar):

45% of the number =86 45% * n = 86 [ we have taken the number is n]

OpenStudy (flygirlmandi):

so 86 divided by 0.45 is n (im only in 7th grade)

OpenStudy (dayakar):

45% *n = 86 0.45n = 86 [ 45% = 0.45] is it clear then check options u can get the answer

OpenStudy (flygirlmandi):

its c?

OpenStudy (dayakar):

no

OpenStudy (flygirlmandi):

im so confused

OpenStudy (dayakar):

what is our final answer

OpenStudy (flygirlmandi):

b? or a?

OpenStudy (dayakar):

why u are confusing ? it is very simple check again it has only one answer

OpenStudy (flygirlmandi):

oh a its a isn't it

OpenStudy (dayakar):

exactly 0.45n =86 86 = 0.45n

OpenStudy (dayakar):

is it clear

OpenStudy (flygirlmandi):

i think i feel really stupid

OpenStudy (flygirlmandi):

thanks though

OpenStudy (dayakar):

don't feel like that , no body is expert

OpenStudy (flygirlmandi):

:)